OpenStudy (anonymous):

Using quadratic formula -2x^2+x+8=0

OpenStudy (muzzack):

Step by step solution : Step 1 : Simplify \[ x-2x^2+8\]

OpenStudy (muzzack):

Trying to factor by splitting the middle term 1.1 Factoring \[-2x^2+x+8 \]

OpenStudy (muzzack):

The first term is, -2x^2 its coefficient is -2 . The middle term is, +x its coefficient is 1 . The last term, "the constant", is +8

OpenStudy (muzzack):

Step-1 : Multiply the coefficient of the first term by the constant -2 • 8 = -16 Step-2 : Find two factors of -16 whose sum equals the coefficient of the middle term, which is 1 .

OpenStudy (muzzack):

are you understanding this?

OpenStudy (muzzack):

-16 + 1 = -15 -8 + 2 = -6 -4 + 4 = 0 -2 + 8 = 6 -1 + 16 = 15 Observation : No two such factors can be found!! Conclusion : Trinomial can not be factored

OpenStudy (muzzack):

Equation at the end of step 1 : \[ -2x^2+x+8=0 \]

OpenStudy (muzzack):

Step 2 : Solve -2x^2+x+8 = 0

OpenStudy (muzzack):

Solve Quadratic Equation by Completing The Square Solving -2x^2+x+8 = 0 by Completing The Square . Multiply both sides of the equation by (-1) to obtain positive coefficient for the first term: 2x^2-x-8 = 0 Divide both sides of the equation by 2 to have 1 as the coefficient of the first term : x^2-(1/2)x-4 = 0

OpenStudy (muzzack):

Add 4 to both side of the equation : x^2-(1/2)x = 4 Now the clever bit: Take the coefficient of x , which is 1/2 , divide by two, giving 1/4 , and finally square it giving 1/16 Add 1/16 to both sides of the equation : On the right hand side we have : 4 + 1/16 or, (4/1)+(1/16) The common denominator of the two fractions is 16 Adding (64/16)+(1/16) gives 65/16 So adding to both sides we finally get : x^2-(1/2)x+(1/16) = 65/16 Adding 1/16 has completed the left hand side into a perfect square : x^2-(1/2)x+(1/16) = (x-(1/4)) • (x-(1/4)) = (x-(1/4))2 Things which are equal to the same thing are also equal to one another. Since x^2-(1/2)x+(1/16) = 65/16 and x^2-(1/2)x+(1/16) = (x-(1/4))2 then, according to the law of transitivity, (x-(1/4))2 = 65/16 We'll refer to this Equation as Eq. #2.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (x-(1/4))2 is (x-(1/4))2/2 = (x-(1/4))1 = x-(1/4) Now, applying the Square Root Principle to Eq. #2.2.1 we get: x-(1/4) = √ 65/16 Add 1/4 to both sides to obtain: x = 1/4 + √ 65/16 Since a square root has two values, one positive and the other negative x^2 - (1/2)x - 4 = 0 has two solutions: x = 1/4 + √ 65/16 or x = 1/4 - √ 65/16 Note that √ 65/16 can be written as √ 65 / √ 16 which is √ 65 / 4 Solve Quadratic Equation using the Quadratic Formula 2.3 Solving -2x2+x+8 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = -2 B = 1 C = 8 Accordingly, B2 - 4AC = 1 - (-64) = 65 Applying the quadratic formula : -1 ± √ 65 x = ————— -4 √ 65 , rounded to 4 decimal digits, is 8.0623 So now we are looking at: x = ( -1 ± 8.062 ) / -4 Two real solutions: x =(-1+√65)/-4=-1.766 or: x =(-1-√65)/-4= 2.266 Two solutions were found : x =(-1-√65)/-4= 2.266 x =(-1+√65)/-4=-1.766